Calculate the pH and pOH of each of the following solutions.

1) [H3O+]= 1.9×10−8 M
2) [H3O+]= 1.0×10−7 M
3) [H3O+]= 2.3×10−6 M
1
Expert's answer
2020-04-02T09:03:27-0400

"pH=-log_{10}[H^{+}]"

"pH+pOH=14"


1) "pH= -log_{10}[1.9\u00d710^{-8}]=7.7"

"pOH=14-7.7= 6.3"


2) "pH= -log_{10}[1.0\u00d710^{-7}]=7"

"pOH=14-pOH=14-7=7"


3) "pH=-log_{10}[2.3\u00d710^{-6}]=5.6"

"pOH=14-5.6=8.4"